The sum of two consecutive odd address numbers on a street is 9044. What are the two address numbers? You must construct an algebra equation to solve this
Answer by checkley77(12844)

x+(x+2)=9044
2x=9044-2
2x=9042
x=9042/2
x=4521 ans for the lower numbered address.
4521+2=4523 ans. for the larger numbered address.
Proof:
4521+4523=9044
9044=9044